------------------------------------------------------------ Date: Sat Aug 25 08:19:58 2001, Shift: 00:00-08:00 Supervisor: J.I.Jordre, with: Log book#8 pages 18 - 20 Shift Summary: ------------------------------------------------------------ 1. Stable physics running beam, except for 2.5 hours qhench link interlock and recovery from 1:20 to 3:50. 2. (Since 8/18: Magnets @ A polarity, D1-D4 @ 1/3 field and D5 @ 2kG.) Run 4395 (from last shift) with 86.5k trigger 1 events. Swapped to D1 @ 1/5 field and other magnets @ 0 field. Run 4399 with 123.3 k trigger 1 events. Swapped to B polarity, D1-D4 @ 1/5 field and D5 @ 2kG. Run 4400 still running. 3. None ------------------------------------------------------------
This archive was generated by hypermail 2b30 : Sat Aug 25 2001 - 08:20:37 EDT